### v1

# BugMagnifier: TON Transaction Simulator for Revealing Smart Contract Vulnerabilities

Abstract
The Open Network (TON) blockchain employs an asynchronous execution model that introduces unique security challenges for smart contracts, particularly race conditions arising from unpredictable message processing order. While previous work established vulnerability patterns through static analysis of audit reports, dynamic detection of temporal dependencies through systematic testing remains an open problem. We present BugMagnifier, a transaction simulation framework that systematically reveals vulnerabilities in TON smart contracts through controlled message orchestration. Built atop TON Sandbox and integrated with the TON Virtual Machine (TVM), our tool combines precise message queue manipulation with differential state analysis and probabilistic permutation testing to detect asynchronous execution flaws. Experimental evaluation demonstrates BugMagnifier’s effectiveness through extensive parametric studies on purpose-built vulnerable contracts, revealing message ratio-dependent detection complexity that aligns with theoretical predictions. This quantitative model enables predictive vulnerability assessment while shifting discovery from manual expert analysis to automated evidence generation. By providing reproducible test scenarios for temporal vulnerabilities, BugMagnifier addresses a critical gap in the TON security tooling, offering practical support for safer smart contract development in asynchronous blockchain environments.
publicationid: pubid: 0000, 0000/00\$00.00 © 2021 IEEE
I Introduction
Blockchain architectures face an inherent tension between transactional atomicity and network scalability, a fundamental trade-off defining platform security profiles [1, 2, 3]. Ethereum’s synchronous model [4] demonstrates this through complex atomic operations such as decentralized exchange (DEX) arbitrage [5, 6], enabling coordinated asset swaps within single transactions. These operations rely on deterministic execution ordering, ensuring static market conditions during processing. However, this atomicity restricts throughput to approximately 30 transactions per second (TPS), limiting scalability to single-chain performance.
The Open Network (TON) addresses scalability through a sharded architecture where transactions decompose into independent messages processed concurrently across chains [7, 8]. This asynchronous model enables theoretical throughput exceeding 100,000 TPS, i.e., an improvement of over three orders of magnitude. However, this innovation introduces novel security challenges. Where Ethereum implements arbitrage atomically, TON requires decomposition into interdependent messages. Developers must manage state changes between executions, cross-shard fees, and partial failure recovery. Unpredictable message interleaving creates distinct vulnerability patterns, with temporal dependencies becoming predominant.
TON’s architecture inherently amplifies race condition risks through five-phase transaction processing , where intermediate states become visible to subsequent messages. The cell-based storage of the TVM exacerbates risks through dynamic fee calculations based on the depth of the cell tree, enabling financial attacks via strategic message sequences [9]. Official TON security guidelines explicitly warn about race conditions in message flows and account destruction scenarios, categorizing them as critical vulnerabilities requiring preventive measures.
While Ethereum’s ecosystem has experienced critical vulnerabilities like The DAO’s reentrancy attack and Parity’s access control flaws [10, 11, 12], these differ fundamentally from TON’s architecturally induced race conditions. Ethereum’s documented race condition vulnerabilities [13] typically require deliberate transaction ordering rather than emerging from inherent design constraints. TON’s immature security ecosystem compounds these challenges; existing analyzers like TONScanner [14] and TSA [15] focus on code defects rather than temporal vulnerabilities, leaving developers without systematic methods to detect message-ordering flaws.
This paper addresses this critical need by introducing BugMagnifier, a transaction scheduling and simulation framework designed specifically to reveal vulnerabilities in TON smart contracts through controlled message orchestration. Unlike conventional analysis tools, BugMagnifier operates at the message granularity level, enabling precise manipulation of message queues and TVM states to systematically expose temporal dependencies. The framework implements a novel permutation testing methodology that quantifies vulnerability manifestation probability based on message ratios, addressing the core challenge of non-deterministic message ordering in TON’s architecture. The key contributions of this paper are as follows:
  1. 1.
    Designed and developed the TON-Complete Transaction Simulation that integrates with TON Sandbox to enable local emulation of TON’s complete transaction processing cycle, including all five execution phases, a capability absent in existing tools that typically support only the compute phase.
  2. 2.
    Implemented a message permutation engine that systematically explores execution paths through configurable message ordering, enabling detection of race conditions that would remain undetected by conventional testing methodologies. The engine incorporates probabilistic analysis to identify vulnerable message ratios and quantify detection complexity.
  3. 3.
    Established a quantitative model of a race condition manifestation in TON smart contracts, validated through extensive execution permutations on a purpose-built vulnerable contract. Our experiments reveal a message ratio-dependent detection complexity that aligns with theoretical predictions, providing actionable insights for security testing protocols.
By shifting vulnerability discovery from manual expert analysis to systematic evidence generation, BugMagnifier addresses a critical gap in TON security tooling. The framework provides reproducible test scenarios for asynchronous execution flaws, particularly race conditions, that have proven exceptionally challenging to detect through conventional methods. Our work contributes to a more secure and robust blockchain environment by empowering developers with practical tools to build safer and more reliable decentralized applications on TON, ultimately supporting the ecosystem’s maturation and adoption.
The remainder of this paper is organized as follows. Section II provides a technical foundation of TON’s architecture with a specific focus on the security implications of its asynchronous execution model. Section III analyzes the limitations of existing security tools in addressing TON’s unique vulnerability landscape. Section IV details BugMagnifier’s design, implementation, and technical innovations. Section V demonstrates the capabilities of the framework through a deliberately vulnerable smart contract. Section VI presents an empirical validation of our approach through comprehensive numerical experiments. Finally, Section VII discusses implications for TON security practices and future research directions.
II Background
II-A TON Virtual Machine (TVM)
The TON Virtual Machine (TVM) serves as the deterministic execution environment for smart contracts on The Open Network, operating as a stack-based virtual machine with specific architectural features tailored to TON’s sharded blockchain design [16]. As blockchains operate in an untrusted environment where participants must self-verify computational consistency, computation results need to be hardware-independent; virtual machines like TVM are the standard choice to ensure this determinism across all network nodes. Unlike Ethereum’s virtual machine (EVM), which processes transactions as atomic units, TVM is specifically engineered to handle TON’s message-centric execution model, where transactions decompose into independent messages processed across (parallel) shard chains. The TVM state comprises five critical components: Stack (for data manipulation), Control Registers (managing execution flow), Current Continuation (tracking program state), Current Codepage (storing executable code), and Gas Limits (enforcing computational constraints).
To prevent deadlock and ensure liveness, the user program execution in TVM, as in EVM, is strictly limited in both computation and memory. This limitation is expressed in universal gas units. The protocol pre-defines a cost in gas for every elementary operation and for storage, forming the basis for transaction fees that users are charged based on the gas consumed by their transactions.
TVM implements a comprehensive instruction set optimized for blockchain operations, featuring 256 primitive opcodes organized into logical groups. The stack manipulation primitives include basic operations like PUSH, POP, and XCHG, as well as compound operations such as PUXC and XCHG3 that enhance code density. For data handling, TVM utilizes cells as its fundamental data structure, recursive tree-like containers that store serialized data in bit strings and references to other cells. This cell-based architecture enables efficient Merkle tree constructions essential for TON’s sharding mechanism. TVM’s cell operations include specialized primitives like STORING (for storage fee calculation) and ACCEPT_MSG_VALUE (for credit phase processing), which directly impact security considerations in smart contract development.
The TVM instruction set incorporates blockchain-specific functionality through application-oriented primitives, including message handling (SEND_RAW_MESSAGE), gas management (ACCEPT_MSG_VALUE), and cryptographic operations (SHA256_C). Notably, TVM implements a five-phase transaction processing model: storage, credit, compute, action, and bounce, each with distinct gas accounting rules and state transition semantics. This multi-phase execution model creates unique security challenges, as vulnerabilities can manifest across phase boundaries rather than within isolated transaction contexts. The compute phase, where contract logic is executed, operates under strict gas limits that vary based on preceding phase outcomes, requiring precise gas calculation to avoid out-of-gas failures that could leave contracts in inconsistent states.
This phased gas model within TVM is a direct consequence of, and is designed to support, TON’s larger asynchronous message-passing architecture.
II-B Asynchronous Transaction Execution
TON’s asynchronous execution model fundamentally differs from conventional blockchains through its message-driven architecture, where transactions decompose into independent messages processed concurrently across shard chains. Each message undergoes five sequential processing phases: storage (calculating storage fees), credit (accounting for incoming funds), compute (executing contract logic), action (generating new messages), and bounce (handling failed message processing). This phased execution enables higher throughput by allowing parallel processing of independent messages while maintaining deterministic outcomes through the consensus algorithm that establishes a final total order for messages.
The asynchronous model introduces critical security considerations absent in synchronous blockchains. While the consensus algorithm ensures a final deterministic outcome across the network, the concurrent processing of messages across shards creates a period of temporal non-determinism from the perspective of a single smart contract. The order in which messages from different senders interleave in their queue is not fixed until included in a block, creating race conditions when multiple messages target the same contract state. Unlike Ethereum’s transaction atomicity, where state changes occur only upon successful transaction completion, TON contracts experience intermediate state transitions after each message processing phase, particularly during the compute phase, which can be observed by subsequent messages. This behavior enables temporal vulnerabilities where the contract state appears consistent during individual message processing but becomes inconsistent when messages interleave unexpectedly.
The bounce phase represents a unique TON security consideration, where failed message processing triggers automatic fund recovery through bounce messages. Contracts must implement proper bounce message handling using the SET_BOUNCE primitive; failure to do so can result in permanent fund loss during message processing failures. Additionally, TON’s storage fee mechanism, calculated dynamically during the storage phase based on cell tree depth, creates financial attack vectors where malicious actors could force contracts into storage debt through carefully crafted message sequences.
II-C Comparative Analysis of Execution Models
The architectural divergence between TON and Ethereum creates significant conceptual challenge for developers transitioning between ecosystems. In Ethereum’s synchronous model, transactions execute atomically within blocks, ensuring that intermediate states remain invisible to external contracts until finalization. This guarantees temporal consistency where state changes occur only upon successful transaction completion. Conversely, TON’s asynchronous model exposes intermediate contract states after each message processing phase, requiring developers to design contracts that maintain consistency across partial execution states.
The most critical translation challenge involves concurrency handling: Ethereum developers typically assume sequential transaction execution, while TON developers must explicitly account for message interleaving. For instance, an Ethereum contract implementing a simple token transfer would execute as a single atomic operation, whereas the equivalent TON implementation requires careful state management across multiple messages, deposit, validation, and withdrawal, each potentially interleaved with messages from other users. This fragmentation necessitates explicit synchronization mechanisms, such as nonce-based state validation or message sequencing constraints, which have no direct Ethereum equivalents.
Another fundamental difference lies in error handling semantics. Ethereum transactions revert entirely upon failure, while TON’s phased execution allows partial state changes to persist even when subsequent phases fail. The bounce phase specifically handles message processing failures by generating recovery messages, requiring contracts to implement defensive patterns that prevent state inconsistencies when bounce messages interleave with new operations. These architectural differences necessitate a complete rethinking of security patterns, as conventional Ethereum security practices like reentrancy guards become insufficient for addressing TON’s message interleaving vulnerabilities.
II-D Implications for Smart Contract Development
The asynchronous execution model of TON necessitates fundamentally different development practices compared to synchronous blockchains. Developers must implement explicit state validation at each message processing stage, as contracts can be invoked with partially updated states from previous incomplete transactions. This requirement transforms seemingly simple operations, like balance transfers, into complex state machines where intermediate states must remain consistent under arbitrary message interleaving.
Security-critical contracts require sophisticated message sequencing constraints to prevent temporal vulnerabilities. For example, deposit-withdrawal patterns must incorporate nonce mechanisms or timestamp validation to ensure message processing order integrity, as the natural message queue ordering provides no guarantees of sequential execution. The storage fee mechanism further complicates development, as contracts must maintain sufficient balance not only for operational gas but also for dynamic storage costs that vary with cell tree depth, a vulnerability vector where attackers can force contracts into storage debt through strategically crafted messages.
TON’s bounce message system introduces additional security considerations, as contracts must properly handle both successful and bounced message processing paths. Failure to implement comprehensive bounce message validation can lead to fund loss or state corruption when processing interleaved message sequences. The phased transaction model also requires precise gas budgeting across all five phases, as gas exhaustion at any stage produces different failure modes, from partial state updates to complete transaction rollback, depending on the failure phase.
These complexities necessitate specialized security testing approaches that explicitly model message ordering permutations, as conventional test methodologies designed for atomic transactions cannot adequately validate contract behavior under TON’s asynchronous execution model. The lack of temporal determinism in message processing requires developers to adopt formal verification techniques or systematic permutation testing to ensure contract resilience against race conditions and other temporal vulnerabilities.
III Related Work
Unlike the well-studied Ethereum ecosystem, smart contract security in TON remains an under-researched area, primarily due to the platform’s architectural divergence from conventional blockchain models. TON’s asynchronous execution model, sharded architecture, and cell-based data structures introduce unique security challenges that render many Ethereum-focused analysis techniques ineffective. The fundamental difference lies in TON’s message-driven execution paradigm, where transactions decompose into independent messages processed across parallel chains, creating vulnerability patterns fundamentally distinct from those observed in synchronous blockchains.
To address this gap, Song et al. [14] conducted a pioneering study identifying eight key vulnerability types specific to TON, Bad Randomness, Precision Loss, Unchecked Return, Global Var Redefined, Improper Function Modifier, Unchecked Bounced Message, Inconsistent Data, Lack End Parse, and proposed TONScanner, a static analysis framework designed to detect these defects. TONScanner leverages the FunC compiler’s frontend to transform FunC source code into an intermediate representation (IR) structured as a directed acyclic graph (DAG). The tool then constructs a control flow graph (CFG) and converts it into static single assignment (SSA) form to streamline analysis. To accommodate TON’s unique data structures, TONScanner integrates specialized components such as Data Dependency Analysis, Call Graph Construction, Taint Analysis, and Cell Construct Analysis. These modules collectively enable the detection of the eight identified defect categories.
Technically, TONScanner’s Cell Construct Analysis module addresses TON’s distinctive cell-based storage mechanism by implementing a recursive cell tree validator that checks for proper cell serialization and deserialization patterns. This component specifically targets vulnerabilities arising from improper cell handling, such as cell overflow conditions and incorrect bit string manipulations, issues that are nonexistent in Ethereum’s account-based model. The tool’s Taint Analysis engine has been extended with TON-specific sinks, including message sending operations (SEND_RAW_MESSAGE) and bounce handling mechanisms (SET_BOUNCE), which are critical for identifying message redirection vulnerabilities.
In an evaluation of 1,640 real-world TON smart contracts, TONScanner uncovered 14,995 defects, highlighting the prevalence of developer errors in the ecosystem. A manual audit of randomly sampled results confirmed the tool’s 97.49% precision, demonstrating its reliability for practical use. Notably, the study revealed that Unchecked Bounced Message vulnerabilities accounted for 38.7% of all detected issues, underscoring the significance of TON’s unique bounce phase in security analysis, a vulnerability class absent in Ethereum ecosystems.
A more recent advancement in TON smart contract analysis is the Ton Symbolic Analyzer (TSA) [15], which operates at the bytecode level using symbolic virtual machines and satisfiability modulo theories (SMT) solvers. This tool is specifically designed to: (1) detect potential runtime errors in the TON Virtual Machine (TVM), (2) generate regression tests based on Blueprint framework specifications, and (3) identify malicious honeypot contracts, a class of deceptive smart contracts engineered to exploit users. TSA demonstrates notable effectiveness in flagging fraudulent contract behavior, verifying expected program execution paths, and detecting anomalous runtime conditions.
TSA’s technical innovation lies in its TVM-aware symbolic execution engine, which models TON’s five-phase transaction processing (storage, credit, compute, action, and bounce) as distinct symbolic states. This approach enables the tool to detect vulnerabilities that span multiple transaction phases, particularly critical for identifying race conditions that manifest across message boundaries. The implementation incorporates a custom SMT theory for TON’s cell operations, allowing precise modeling of cell serialization vulnerabilities that would be undetectable through conventional symbolic execution. The evaluation [15] demonstrated TSA’s capability to identify complex multi-phase vulnerabilities with 89.3% recall, though at significant computational cost (average 27 minutes per contract).
While existing tools represent important initial efforts in addressing TON-specific vulnerabilities, they lack the comprehensive coverage and maturity of their Ethereum counterparts. As demonstrated in our previous work [17], which established a detailed taxonomy of TON smart contract vulnerabilities and proposed a structured audit methodology, there remains a significant gap in practical analysis tools for this emerging ecosystem. Most critically, current solutions fail to address the temporal dimension of TON security, specifically, the ability to systematically explore message ordering permutations and their impact on contract state. Static analysis tools such as TONScanner cannot model runtime message interleaving, while symbolic execution approaches like TSA become computationally intractable when analyzing multiple message sequences.
In the present study, we introduce BugMagnifier, an advanced debugging instrument designed specifically for TON smart contract analysis. This tool complements our earlier theoretical framework by providing: (1) dynamic analysis capabilities through precise message queue manipulation, (2) enhanced vulnerability detection mechanisms for temporal vulnerabilities via message order permutations, and (3) real-time debugging features tailored to TON’s unique architecture, including full transaction phase visibility and cell-level state inspection. Unlike existing solutions, BugMagnifier operates at the message granularity level, enabling systematic exploration of execution paths that span multiple transactions, a capability essential for identifying the temporal vulnerabilities that dominate TON’s security landscape.
Through this combined theoretical and practical approach, we aim to bridge the gap between Ethereum’s mature audit methodologies and the evolving needs of the TON ecosystem. BugMagnifier specifically addresses the critical deficiency in current tooling by providing the first practical solution for systematic race condition detection in TON smart contracts, a vulnerability class responsible for the majority of critical security incidents in the ecosystem.
IV Proposed Solution
This section provides a comprehensive description of BugMagnifier, a specialized framework designed to detect and demonstrate vulnerabilities in smart contracts within the asynchronous transaction model of the TON blockchain. Unlike conventional blockchain analysis tools that operate at the transaction level, BugMagnifier introduces message-level execution control, a critical capability for exposing temporal vulnerabilities that manifest only under specific message interleaving sequences. The framework’s architecture centers on precise manipulation of message processing order while maintaining complete visibility into contract state transitions, enabling systematic exploration of execution paths that would remain hidden in conventional testing approaches. This methodology addresses the fundamental challenge of non-deterministic message ordering in the TON’s architecture, where vulnerability manifestation depends on the sequence of message processing rather than isolated transaction properties.
IV-A Choosing a Blockchain Emulator
The foundation of BugMagnifier’s effectiveness lies in its ability to model TON’s complete transaction processing mechanics with surgical precision. To achieve this, the emulator must satisfy three non-negotiable requirements derived from TON’s architectural constraints:
  1. 1.
    comprehensive support for all five transaction phases (storage, credit, compute, action, and bounce),
  2. 2.
    access to intermediate TVM states between phase transitions,
  3. 3.
    extensibility to manipulate message queue ordering.
These requirements emerged from the recognition that temporal vulnerabilities in TON contracts often manifest across phase boundaries rather than within isolated execution contexts.
Our evaluation revealed critical limitations in existing tools. TON Contract Executor [18] abstracts away all phases except compute, bypassing essential TVM opcodes such as ACCEPT_MSG_VALUE and STORING. This creates dangerous blind spots for vulnerabilities involving storage fees or bounce message handling. Similarly, TVM-linker [19] implements storage fees as fixed overhead rather than through TON’s dynamic cell depth calculations and omits bounce phase processing entirely. Both tools fail to preserve intermediate TVM states between phase transitions, rendering them incapable of detecting state inconsistencies that occur during partial execution sequences.
TON Sandbox [20] emerged as the only viable foundation due to its transparent implementation of TON’s transaction lifecycle. Its transaction processor module exposes the complete TVM state machine with precise gas accounting at each opcode execution, preserving full context of intermediate states between phase transitions. Crucially, its modular architecture provides hooks into the message processing pipeline, enabling BugMagnifier’s core innovation: a message interceptor that dynamically reorders messages while maintaining visibility into all five transaction phases. This capability allows systematic exploration of message sequences that trigger temporal vulnerabilities, particularly race conditions where contract behavior depends on the interleaving of messages from multiple senders.
Building on TON Sandbox, BugMagnifier implements three critical enhancements: (1) a state snapshotting mechanism capturing complete TVM register states at each execution phase with minimal overhead (12%); (2) a differential analyzer comparing memory layouts at the cell level to detect subtle inconsistencies; and (3) a message timing simulator modeling network latency effects on message interleaving. These innovations enable precise reproduction of edge cases where vulnerabilities manifest only under specific network conditions, such as storage debt scenarios triggered by strategically crafted cell tree depths or race conditions dependent on message arrival timing.
When initialized, BugMagnifier creates a local blockchain instance with capabilities exceeding standard development environments. The framework implements a copy-on-write state preservation system allowing rollback to any execution point, granular gas consumption control to simulate partial phase completions, and comprehensive execution tracing that maps message sequences to specific state transitions across all transaction phases. These features establish BugMagnifier as the first practical solution for systematic detection of temporal vulnerabilities in TON smart contracts, a class of flaws yet undetectable by conventional analysis methods.
IV-B BugMagnifier
The BugMagnifier interactive console is a powerful tool that provides developers with the ability to emulate TON transactions locally and identify vulnerabilities in smart contracts through advanced message management and simulation of various execution scenarios. The tool is publicly available on GitHub [21]. Built as a TypeScript application leveraging Node.js runtime, BugMagnifier implements a layered architecture with distinct components for contract compilation, transaction simulation, state management, and analysis. At its core, the tool establishes a bidirectional communication channel with the TON Sandbox through a custom IPC (Inter-Process Communication) protocol that enables fine-grained control over transaction execution while maintaining compatibility with the TVM’s deterministic execution requirements.
During initialization, users specify the path to the target smart contract. Behind the scenes, BugMagnifier compiles the source code into executable TVM bytecode, saving the compiled code as a JSON file with base64 cell representation in the temporary directory tmp/tondebug.compiled.json. The compilation process utilizes the official FunC compiler with customized error reporting that maps bytecode offsets to source code locations, enabling precise debugging of complex contracts. Users can optionally configure the initial smart contract state and message queue before console launch, though both remain editable during interactive sessions. The initialization sequence implements a verification step that checks the contract’s compliance with TON’s compute phase limits, pre-emptively identifying potential out-of-gas conditions before execution begins.
Smart contract state configuration uses a JSON structure with three fields: contract balance, hex representation of TVM code, and base64-serialized contract state data. The balance parameter is particularly useful for simulating financial scenarios like insufficient balance handling during message processing or storage payments. Example state configuration:
1{
2 "balance": "1099088800",
3 "data": "b5ee9c...240000438014ffc...",
4 "code": "b5ee9c...30000114Ff00f4a..."
5}
The message queue employs a JSON array where each message contains these fields: unique identifier id, auto-generated if unspecified, internal/external-in type, defaults to internal, payload body interpreted by contract logic, nanotoken transfer amount value for internal messages, numerical sender identifier senderId, random if omitted, and optional test case label name. Example message queue:
1[
2 {
3 "id": 1,
4 "type": "internal",
5 "body": "te6ccgEBAQEABgAACAAAAAI=",
6 "value": {"coins": "10000000"},
7 "senderId": 1,
8 "name": "CLAIM Alice"
9 },
10 {
11 "id": 2,
12 "type": "internal",
13 "body": "te6ccgEBAQEABgAACAAAAAE=",
14 "value": {"coins": "10000000"},
15 "senderId": 2,
16 "name": "ENLIST Bob"
17 }
Execution begins with the command:
1tondebug , contract <path> [, init-state <path>]
2[, queue <path>] [, help]
The console provides comprehensive message queue control through commands in the table I.
TABLE I: BugMagnifier Command Reference
Command Functionality
run next Processes next message in queue, recording final state, transaction details, and message content
run message <id> Executes message by ID (errors on invalid IDs)
continue Processes all remaining messages
queue list Displays current message queue
set queue , order Modifies execution order (reverse or random)
add messages <path> Adds new messages via JSON file
delete message <id> Removes message by ID
script load <path> Loads TS/JS script with modifyQueue()
script run Applies loaded script to queue
show state Displays contract status
show transactions Shows transaction log
show message log Displays message history
load state <path> Loads new contract state
save state <path> Saves current state to JSON
diff <path1> <path2> Compares contract states
exit Terminates console
help Displays documentation
Scripting examples demonstrate queue modification capabilities:
JavaScript randomization script:
1/**
2@param {import(’../tondebug/tondebug.ts’)
3.Message[]} queue
4*/
5export function modifyQueue(queue) {
6for (let i = queue.length - 1;
7i > 0; i, ) {
8const j = Math.floor(
9Math.random() * (i + 1));
10[queue[i], queue[j]] =
11[queue[j], queue[i]];
12}
13}
This script randomizes the message execution order using the Fisher-Yates shuffle algorithm to simulate non-deterministic race conditions and uncover ordering-dependent vulnerabilities.
TypeScript prioritization script:
1export function modifyQueue(
2queue: Message[]): void {
3queue.sort((a, b) => {
4const byType =
5typePriority[a.type] -
6typePriority[b.type];
7if (byType !== 0) return byType;
8const coinsA = a.value?.coins ?? 0n;
9const coinsB = b.value?.coins ?? 0n;
10return coinsB > coinsA ? 1 : coinsB < coinsA ? -1 : 0;
11});
12}
This script prioritizes messages by first comparing their type (e.g., external vs. internal) and then by descending value (coins), enabling testing of scenarios where message financial value or type determines processing priority.
The scripting interface implements a secure sandbox environment that restricts access to system resources while providing full access to the message queue structure. This design enables complex test scenarios without compromising system security. Advanced users can implement sophisticated vulnerability detection algorithms through the scripting API, including probabilistic analysis of state transitions and custom anomaly detection heuristics.
The tool provides comprehensive control over smart contract testing by combining precise execution management (launching, removing, or adding specific messages) with automation capabilities (processing entire message queues) and thorough state analysis. This system enables: (1) processing order control through configurable message queues and scriptable race condition simulations; (2) change visualization via execution history and queue monitoring; and (3) systematic recording and analysis of all contract state changes post-transaction, a critical capability for vulnerability detection that includes simulating complex attacks and assessing their impact on contract state. The differential analysis engine implements a cell-level comparison algorithm that identifies subtle state inconsistencies that are invisible to higher-level inspections, such as differences in cell bit patterns or hash tree structures. This capability is particularly valuable for detecting vulnerabilities involving complex data structures where surface-level state comparisons might miss critical discrepancies.
The solution supports the complete testing lifecycle, from environment initialization (automated FunC to TVM bytecode compilation and JSON-configured state setup) to post-execution analysis (including state comparison and race condition detection). Performance optimizations include a state caching mechanism that reduces redundant computation during repeated executions, and a parallel execution mode that leverages multi-core processors to accelerate permutation testing. Of particular value is the interactive mode, which allows dynamic message queue manipulation, execution order modification (sequential, reverse, or randomized), and integration of custom scripts for specialized test scenarios. The tool’s architecture implements a plugin system that enables extension of core functionality without modifying the base code, facilitating integration with external analysis tools and continuous integration pipelines.
This combination of features establishes BugMagnifier as an essential tool for verifying smart contract’s reliability and security prior to deployment on the live TON blockchain network. The ability to precisely control execution parameters and capture complete state transitions makes BugMagnifier particularly effective at identifying temporal vulnerabilities that would be nearly impossible to detect through conventional testing methods.
V Illustrative Vulnerable Smart Contract
As part of the illustrative numerical experiment, we employ the BugMagnifier tool to demonstrate the complexities of detecting race conditions in TON smart contracts. To facilitate this analysis, we designed a deliberately vulnerable FunC smart contract implementing a deposit pool mechanism that exhibits critical race conditions under concurrent message processing. This example reveals how subtle implementation details in asynchronous blockchain environments can lead to severe financial inconsistencies, even in seemingly straightforward contracts. The vulnerability stems from improper state management during the ownership assignment phase, a critical weakness that becomes apparent only when analyzing message sequence permutations.
The test contract implements a deposit pool mechanism centered around two core state variables: a 64-bit integer representing the accumulated balance (balance) and a slice storing the owner’s address (owner_address). Upon initial deployment, the contract establishes a zero balance and a null owner address (represented by an empty slice). Transaction processing logic is implemented through two primary operations: OP_ENLIST (value 1) for balance deposits and OP_CLAIM (value 2) for withdrawals. The critical vulnerability emerges from the contract’s ownership assignment mechanism, which triggers during the first deposit operation. This implementation pattern, while seemingly logical for single-user initialization, creates a fundamental race condition when multiple deposit messages arrive concurrently.
The following listing presents the complete vulnerable contract implementation with critical sections annotated for clarity:
1#include "imports/stdlib.fc";
2
3const OP_ENLIST = 1;
4const OP_CLAIM = 2;
5
6(int, slice) _load_state() {
7 slice ds = get_data()
8 .begin_parse();
9 int balance = slice_bits(ds) >= 64 ?
10 ds.load_uint(64) : 0;
11 slice owner_address = ds;
12 return (balance, owner_address);
13}
14
15(slice) get_msg_sender(cell in_msg) {
16 slice cs = in_msg.begin_parse();
17 int flags = cs.load_uint(4);
18 slice sender_address =
19 cs^.load_msg_addr();
20 return sender_address;
21}
22
23() _save_state(int balance,
24 slice owner_address) impure {
25 set_data(begin_cell()
26 .store_uint(balance, 64)
27 .store_slice(owner_address)
28 .end_cell()
29 );
30}
31
32() recv_internal(int msg_value,
33 cell in_msg, slice in_msg_body)
34 impure {
35 var (balance, owner_address) =
36 _load_state();
37 slice sender_address =
38 get_msg_sender(in_msg);
39
40 if (slice_bits(in_msg_body) >= 32) {
41 int op = in_msg_body
42 .load_uint(32);
43 if (op == OP_ENLIST) {
44 balance += msg_value;
45 // CRITICAL VULNERABILITY POINT
46 // Ownership assignment occurs ONLY
47 // during first deposit when
48 // owner_address is empty (0 bits)
49 if (slice_bits(owner_address) == 0) {
50 _save_state(balance,
51 sender_address);
52 } else {
53 _save_state(balance,
54 owner_address);
55 }
56 return();
57 }
58 if (op == OP_CLAIM) {
59 // Ownership verification
60 // before withdrawal
61 if (equal_slices(owner_address,
62 sender_address)) {
63 balance = 0;
64 _save_state(balance, null());
65 }
66 return();
67 }
68 }
69}
70
71() recv_external(slice _) {
72 return ();
73}
74
75(int, slice) get_state() method_id {
76 return _load_state();
77}
The vulnerability manifests through non-deterministic message ordering during concurrent execution, a direct consequence of TON’s asynchronous architecture. Consider the following scenario involving three messages:
  1. 1.
    OP_ENLIST from Alice (sender ID 1)
  2. 2.
    OP_ENLIST from Bob (sender ID 2)
  3. 3.
    OP_CLAIM from Alice (sender ID 1).
When these messages arrive concurrently, TON’s message queue processing creates two distinct execution paths with divergent outcomes:
Scenario 1: Bob’s message processes first
  1. 1.
    Bob’s OP_ENLIST executes first
  2. 2.
    Contract detects empty owner_address (0 bits)
  3. 3.
    Bob’s address becomes permanent owner
  4. 4.
    Alice’s subsequent OP_CLAIM fails ownership verification
  5. 5.
    Final state: balance = 20,000,000 nanotokens, owner = Bob.
Scenario 2: Alice’s message processes first
  1. 1.
    Alice’s OP_ENLIST executes first
  2. 2.
    Contract detects empty owner_address
  3. 3.
    Alice’s address becomes permanent owner
  4. 4.
    Alice’s OP_CLAIM succeeds (ownership verified)
  5. 5.
    Final state: balance = 0, owner = null.
This race condition stems from the contract’s dependence on the slice_bits(owner_address) check during ownership assignment, a state transition that should be atomic but is instead exposed to message interleaving. The vulnerability is particularly insidious because:
  • It passes standard functional tests with sequential message processing.
  • It only manifests under specific message order permutations.
  • The financial impact (complete loss of funds) occurs without explicit error conditions.
  • It exploits TON’s asynchronous execution model where message processing order is non-deterministic.
From a technical perspective, the vulnerability arises during the compute phase of transaction execution when the TVM processes the conditional ownership assignment. The critical instruction sequence occurs at the bytecode level as:
1DUP // Duplicate owner_address slice
2SLICELEN // Calculate bit length
3PUSHINT 0 // Compare with zero
4IFEQ // If empty, assign new owner
This sequence creates a race window where concurrent messages can both pass the empty check before either completes state update, a classic check-then-act vulnerability amplified by TON’s message queue architecture.
BugMagnifier enables systematic detection of this vulnerability through its message permutation capabilities. By generating all possible execution sequences of the three-message scenario (6 permutations), the tool identifies two distinct final states:
  1. 1.
    States where Alice becomes owner and withdraws funds (3 permutations)
  2. 2.
    States where Bob becomes owner and funds remain locked (3 permutations).
The differential state analysis reveals critical inconsistencies in both the balance and ownership fields, with the diff command showing:
1State difference detected:
2- balance: 20000000 vs 0
3- owner_address: [Bobs address] vs [empty]
This case study exemplifies how TON’s asynchronous execution model transforms seemingly benign state transitions into critical vulnerabilities. The contract’s failure to implement proper synchronization mechanisms, such as nonce-based ownership claims or explicit state locking, demonstrates the heightened security requirements for asynchronous blockchain development. Our numerical experiments (Section VI) quantify how detection complexity varies with message ratios, revealing that such vulnerabilities become increasingly difficult to identify as message counts grow asymmetrically, a direct consequence of TON’s probabilistic message processing order.
VI Numerical Experiments
The experimental framework generated message queues containing $n_{1}$ messages from Alice and $n_{2}$ messages from Bob, structured as follows:
1[
2 {
3 "id": 1,
4 "type": "internal",
5 "body": "te6ccgEBAQEABgAACAAAAAE=",
6 "value": {"coins": "10000000"},
7 "senderId": 1,
8 "name": "ENLIST Alice (1 TON)"
9 },
10 {
11 "id": 2,
12 "type": "internal",
13 "body": "te6ccgEBAQEABgAACAAAAAE=",
14 "value": {"coins": "10000000"},
15 "senderId": 2,
16 "name": "ENLIST Bob (1 TON)"
17 }
Each configuration was executed using BugMagnifier with the command:
1tondebug , contract contracts/race_head.fc
2, init-state states/initial_rc_state.json
3, queue tmp/generated_queue.json
The experimental protocol executed 1,000 iterations per configuration. In each iteration, the message queue was randomly shuffled and processed sequentially (run message <id>). After the first iteration, the initial state was saved (save state tmp/first_exp_state.json), and subsequent iterations began with a reset environment (load state and add messages). The state divergence between the current execution and baseline, detected through internal state comparison and get-method results (balance and owner address), terminated the iteration sequence.
The methodology evaluated race condition detection complexity across message ratios with fixed Bob messages ($n_{2}=32$) and geometrically increasing Alice messages ($n_{1}=1,2,4,\dots,512$). For each configuration, For each configuration, 100 independent trials recorded the iteration count where state divergence first occurred, providing a quantitative measure of a contract’s susceptibility to race conditions. The experimental results are presented in Table II.
TABLE II: Experimental Results
$n_{1}$ (Alice) $n_{2}$ (Bob) Total Iter. Avg. Iter. Std. Dev.
1 32 2932 29.32 10.45
2 32 2863 28.63 15.08
4 32 932 9.32 8.88
8 32 548 5.48 4.58
16 32 355 3.55 2.58
32 32 306 3.06 1.48
64 32 371 3.71 2.36
128 32 485 4.85 3.87
256 32 972 9.72 12.81
512 32 1546 15.46 12.79
The data reveal a U-shaped relationship between the message ratio and the detection complexity. Minimum detection iterations (3.06) occur at balanced message counts ($n_{1}=n_{2}=32$), while extreme imbalances require significantly more iterations (29.32 for $n_{1}=1$, 15.46 for $n_{1}=512$). Higher standard deviations at imbalanced ratios indicate greater variance in the detection timing.
The theoretical model follows:
$E[X]=\sum_{k=2}^{\infty}k\left(p^{k-1}q+q^{k-1}p\right)$
where $p=\frac{n_{1}}{n_{1}+n_{2}}$ and $q=\frac{n_{2}}{n_{1}+n_{2}}$ represent the relative message probabilities.
Figure 1 plots experimental results against $\log_{2}(n_{1}/n_{2})$ with error bars representing $\pm$1 standard deviation. The x-axis spans from -5 (Bob dominance) to 4 (Alice dominance), with zero representing balanced messages. The experimental data aligns precisely with the theoretical curve , confirming the probabilistic model of race condition manifestation.
Figure 1: Dependence of average detection iterations on message ratio
This correspondence demonstrates that race condition detection complexity follows predictable patterns based on message ratios. Critically, vulnerabilities become substantially harder to detect under highly imbalanced conditions, precisely the scenarios most likely in real-world deployments where certain user groups dominate transaction volume. This finding necessitates targeted testing strategies that account for expected traffic patterns rather than relying on random permutation testing alone.
VII Conclusion
Our research establishes BugMagnifier as a foundational advancement in addressing the unique security challenges posed by TON’s asynchronous execution model. Unlike conventional blockchain security tools designed for atomic transaction processing, BugMagnifier specifically targets the temporal vulnerabilities inherent to TON’s message-driven architecture, particularly race conditions that emerge from non-deterministic message interleaving across the five-phase transaction lifecycle. The framework’s capability to systematically explore message ordering permutations provides the first practical methodology for detecting vulnerabilities that conventional static analysis and symbolic execution approaches cannot reliably identify.
The core innovation lies in BugMagnifier’s message granularity control, which enables precise manipulation of execution flows while maintaining complete visibility into TVM state transitions across all transaction phases. This capability addresses the fundamental limitation of existing tools that typically operate at the transaction level and remain blind to intermediate state changes between message processing stages. By transforming the detection of temporal vulnerabilities from an ad-hoc, expert-dependent process into a systematic, evidence-based methodology, BugMagnifier establishes a new paradigm for security analysis in asynchronous blockchain environments.
The implications of this work extend beyond immediate vulnerability detection. BugMagnifier’s approach establishes a framework for developing security-aware design patterns specifically tailored to TON’s architectural constraints, potentially influencing future smart contract development practices within the ecosystem. The methodology also provides a template for addressing similar challenges in other asynchronous blockchain platforms, suggesting a broader applicability of the core concepts.
Looking forward, this research opens several promising directions. The integration of BugMagnifier’s permutation testing methodology with formal verification techniques could yield hybrid analysis approaches with stronger guarantees. Additionally, the message scheduling capabilities of the framework could inform the development of defensive programming patterns that inherently resist temporal vulnerabilities. As TON continues to mature, tools like BugMagnifier will play a critical role in establishing security standards that match the ecosystem’s technical sophistication, ultimately enabling more robust and reliable decentralized applications on the platform.
Acknowledgments
The research has been supported by the Ministry of Economic Development of the Russian Federation (agreement with MIPT No. 139-15-2025-013, dated June 20, 2025, IGK  [^00] C313925P4B0002).

### v2

@@
 Abstract
-The Open Network (TON) blockchain employs an asynchronous execution model that introduces unique security challenges for smart contracts, particularly race conditions arising from unpredictable message processing order.
+The Open Network (TON) blockchain employs an asynchronous execution model that introduces unique security challenges for smart contracts.
+A primary concern is race conditions arising from unpredictable message processing order.
 While previous work established vulnerability patterns through static analysis of audit reports, dynamic detection of temporal dependencies through systematic testing remains an open problem.
-We present BugMagnifier, a transaction simulation framework that systematically reveals vulnerabilities in TON smart contracts through controlled message orchestration.
-Built atop TON Sandbox and integrated with the TON Virtual Machine (TVM), our tool combines precise message queue manipulation with differential state analysis and probabilistic permutation testing to detect asynchronous execution flaws.
+This study proposes a dynamic evaluation methodology based on controlled message orchestration to systematically expose vulnerabilities in asynchronous smart contracts.
+By synthesizing precise message queue manipulation with differential state analysis and probabilistic permutation testing, we establish a framework (viz, BugMagnifier) for identifying execution flaws that static methods miss.
 Experimental evaluation demonstrates BugMagnifier’s effectiveness through extensive parametric studies on purpose-built vulnerable contracts, revealing message ratio-dependent detection complexity that aligns with theoretical predictions.
@@
 I Introduction
-Blockchain architectures face an inherent tension between transactional atomicity and network scalability, a fundamental trade-off defining platform security profiles [1, 2, 3].
-Ethereum’s synchronous model [4] demonstrates this through complex atomic operations such as decentralized exchange (DEX) arbitrage [5, 6], enabling coordinated asset swaps within single transactions.
+Blockchain architectures face an inherent tension between transactional atomicity and network scalability, a fundamental trade-off defining platform security profiles [Buterin2021a, Kruglik2019a, Amelin2021].
+Ethereum’s synchronous model [Buterin2014] demonstrates this through complex atomic operations such as decentralized exchange (DEX) arbitrage [Gramlich2024, Vostrikov2025], enabling coordinated asset swaps within single transactions.
 These operations rely on deterministic execution ordering, ensuring static market conditions during processing.
 However, this atomicity restricts throughput to approximately 30 transactions per second (TPS), limiting scalability to single-chain performance.
-The Open Network (TON) addresses scalability through a sharded architecture where transactions decompose into independent messages processed concurrently across chains [7, 8].
-This asynchronous model enables theoretical throughput exceeding 100,000 TPS, i.e., an improvement of over three orders of magnitude.
+The Open Network (TON) addresses scalability through a sharded architecture where transactions decompose into independent messages processed concurrently across chains [Durov2019, Berger2019].
+This asynchronous model enables theoretical throughput exceeding 100,000 TPS, i.e.
+an improvement of more than three orders of magnitude.
 However, this innovation introduces novel security challenges.
@@
 TON’s architecture inherently amplifies race condition risks through five-phase transaction processing, where intermediate states become visible to subsequent messages.
-The cell-based storage of the TVM exacerbates risks through dynamic fee calculations based on the depth of the cell tree, enabling financial attacks via strategic message sequences [9].
+Cell-based storage of TVM exacerbates risks through dynamic fee calculations based on the depth of the cell tree, enabling financial attacks via strategic message sequences [TON2025a].
 Official TON security guidelines explicitly warn about race conditions in message flows and account destruction scenarios, categorizing them as critical vulnerabilities requiring preventive measures.
-While Ethereum’s ecosystem has experienced critical vulnerabilities like The DAO’s reentrancy attack and Parity’s access control flaws [10, 11, 12], these differ fundamentally from TON’s architecturally induced race conditions.
-Ethereum’s documented race condition vulnerabilities [13] typically require deliberate transaction ordering rather than emerging from inherent design constraints.
-TON’s immature security ecosystem compounds these challenges; existing analyzers like TONScanner [14] and TSA [15] focus on code defects rather than temporal vulnerabilities, leaving developers without systematic methods to detect message-ordering flaws.
+Whilst Ethereum’s ecosystem has experienced critical vulnerabilities like The DAO’s reentrancy attack and Parity’s access control flaws [Atzei2017, Chen2022a, Soud2024], these differ fundamentally from TON’s architecturally induced race conditions.
+Ethereum’s documented race condition vulnerabilities [Breidenbach2018] typically require deliberate transaction ordering rather than emerging from inherent design constraints.
+The nascent TON security ecosystem currently offers analyzers like TONScanner [Song2025] and TSA [Espirito2025], which are effective for identifying static code defects (e.g., bad randomness, unchecked returns).
+However, these tools operate via static analysis and symbolic execution, methodologies that are inherently unable to simulate the temporal execution paths arising from non-deterministic message interleaving.
+Consequently, a critical gap exists for detecting dynamic, state-dependent race conditions, leaving developers without systematic methods to validate contract behavior under asynchronous execution.
 This paper addresses this critical need by introducing BugMagnifier, a transaction scheduling and simulation framework designed specifically to reveal vulnerabilities in TON smart contracts through controlled message orchestration.
@@
   1. 1.
-    Designed and developed the TON-Complete Transaction Simulation that integrates with TON Sandbox to enable local emulation of TON’s complete transaction processing cycle, including all five execution phases, a capability absent in existing tools that typically support only the compute phase.
+    Designed and developed the TON-Complete Transaction Simulation that integrates with TON Sandbox to enable local emulation of the TON’s complete transaction processing cycle, including all five execution phases, a capability absent in existing tools that typically support only the compute phase.
   2. 2.
@@
 II-A TON Virtual Machine (TVM)
-The TON Virtual Machine (TVM) serves as the deterministic execution environment for smart contracts on The Open Network, operating as a stack-based virtual machine with specific architectural features tailored to TON’s sharded blockchain design [16].
+The TON Virtual Machine (TVM) serves as the deterministic execution environment for smart contracts on The Open Network, operating as a stack-based virtual machine with specific architectural features tailored to TON’s sharded blockchain design [Durov2020].
 As blockchains operate in an untrusted environment where participants must self-verify computational consistency, computation results need to be hardware-independent; virtual machines like TVM are the standard choice to ensure this determinism across all network nodes.
@@
 This multi-phase execution model creates unique security challenges, as vulnerabilities can manifest across phase boundaries rather than within isolated transaction contexts.
-The compute phase, where contract logic is executed, operates under strict gas limits that vary based on preceding phase outcomes, requiring precise gas calculation to avoid out-of-gas failures that could leave contracts in inconsistent states.
+The compute phase, where contract logic is executed, operates under strict gas limits that vary based on preceding phase outcomes, requiring precise gas calculations to avoid out-of-gas failures that could leave contracts in inconsistent states.
 This phased gas model within TVM is a direct consequence of, and is designed to support, TON’s larger asynchronous message-passing architecture.
@@
 TON’s asynchronous execution model fundamentally differs from conventional blockchains through its message-driven architecture, where transactions decompose into independent messages processed concurrently across shard chains.
-Each message undergoes five sequential processing phases: storage (calculating storage fees), credit (accounting for incoming funds), compute (executing contract logic), action (generating new messages), and bounce (handling failed message processing).
+Each message undergoes five sequential processing phases: storage (calculating storage fees), credit (accounting for incoming funds), compute (executing contract logic), action (generating new messages) and bounce (handling failed message processing).
 This phased execution enables higher throughput by allowing parallel processing of independent messages while maintaining deterministic outcomes through the consensus algorithm that establishes a final total order for messages.
-The asynchronous model introduces critical security considerations absent in synchronous blockchains.
+The asynchronous model introduces critical security considerations that are absent in synchronous blockchains.
 While the consensus algorithm ensures a final deterministic outcome across the network, the concurrent processing of messages across shards creates a period of temporal non-determinism from the perspective of a single smart contract.
@@
 Additionally, TON’s storage fee mechanism, calculated dynamically during the storage phase based on cell tree depth, creates financial attack vectors where malicious actors could force contracts into storage debt through carefully crafted message sequences.
+A significant emerging trend is the application of Artificial Intelligence and Large Language Models (LLMs) for automated vulnerability discovery and exploit generation.
+Recent studies demonstrate that models like GPT-4 and Claude can analyze Solidity code and generate functional proof-of-concept exploits for known vulnerability patterns with notable success rates [Xiao2025].
+Although these approaches excel at reasoning about code semantics and predicting flaws based on learned patterns, they operate through static analysis or limited symbolic execution.
+BugMagnifier represents a complementary paradigm: instead of *predicting* vulnerabilities from code, it *concretely executes* contract logic under systematically varied message orderings to *discover* temporal flaws.
+This dynamic exploration-based approach is crucial for detecting asynchronous race conditions that are often poorly represented in the training data of current LLMs.
 II-C Comparative Analysis of Execution Models
@@
 III Related Work
-Unlike the well-studied Ethereum ecosystem, smart contract security in TON remains an under-researched area, primarily due to the platform’s architectural divergence from conventional blockchain models.
-TON’s asynchronous execution model, sharded architecture, and cell-based data structures introduce unique security challenges that render many Ethereum-focused analysis techniques ineffective.
-The fundamental difference lies in TON’s message-driven execution paradigm, where transactions decompose into independent messages processed across parallel chains, creating vulnerability patterns fundamentally distinct from those observed in synchronous blockchains.
-To address this gap, Song et al.
-[14] conducted a pioneering study identifying eight key vulnerability types specific to TON, Bad Randomness, Precision Loss, Unchecked Return, Global Var Redefined, Improper Function Modifier, Unchecked Bounced Message, Inconsistent Data, Lack End Parse, and proposed TONScanner, a static analysis framework designed to detect these defects.
-TONScanner leverages the FunC compiler’s frontend to transform FunC source code into an intermediate representation (IR) structured as a directed acyclic graph (DAG).
-The tool then constructs a control flow graph (CFG) and converts it into static single assignment (SSA) form to streamline analysis.
-To accommodate TON’s unique data structures, TONScanner integrates specialized components such as Data Dependency Analysis, Call Graph Construction, Taint Analysis, and Cell Construct Analysis.
-These modules collectively enable the detection of the eight identified defect categories.
-Technically, TONScanner’s Cell Construct Analysis module addresses TON’s distinctive cell-based storage mechanism by implementing a recursive cell tree validator that checks for proper cell serialization and deserialization patterns.
-This component specifically targets vulnerabilities arising from improper cell handling, such as cell overflow conditions and incorrect bit string manipulations, issues that are nonexistent in Ethereum’s account-based model.
-The tool’s Taint Analysis engine has been extended with TON-specific sinks, including message sending operations (SEND_RAW_MESSAGE) and bounce handling mechanisms (SET_BOUNCE), which are critical for identifying message redirection vulnerabilities.
-In an evaluation of 1,640 real-world TON smart contracts, TONScanner uncovered 14,995 defects, highlighting the prevalence of developer errors in the ecosystem.
-A manual audit of randomly sampled results confirmed the tool’s 97.49% precision, demonstrating its reliability for practical use.
-Notably, the study revealed that Unchecked Bounced Message vulnerabilities accounted for 38.7% of all detected issues, underscoring the significance of TON’s unique bounce phase in security analysis, a vulnerability class absent in Ethereum ecosystems.
-A more recent advancement in TON smart contract analysis is the Ton Symbolic Analyzer (TSA) [15], which operates at the bytecode level using symbolic virtual machines and satisfiability modulo theories (SMT) solvers.
-This tool is specifically designed to: (1) detect potential runtime errors in the TON Virtual Machine (TVM), (2) generate regression tests based on Blueprint framework specifications, and (3) identify malicious honeypot contracts, a class of deceptive smart contracts engineered to exploit users.
-TSA demonstrates notable effectiveness in flagging fraudulent contract behavior, verifying expected program execution paths, and detecting anomalous runtime conditions.
-TSA’s technical innovation lies in its TVM-aware symbolic execution engine, which models TON’s five-phase transaction processing (storage, credit, compute, action, and bounce) as distinct symbolic states.
-This approach enables the tool to detect vulnerabilities that span multiple transaction phases, particularly critical for identifying race conditions that manifest across message boundaries.
-The implementation incorporates a custom SMT theory for TON’s cell operations, allowing precise modeling of cell serialization vulnerabilities that would be undetectable through conventional symbolic execution.
-The evaluation [15] demonstrated TSA’s capability to identify complex multi-phase vulnerabilities with 89.3% recall, though at significant computational cost (average 27 minutes per contract).
-While existing tools represent important initial efforts in addressing TON-specific vulnerabilities, they lack the comprehensive coverage and maturity of their Ethereum counterparts.
-As demonstrated in our previous work [17], which established a detailed taxonomy of TON smart contract vulnerabilities and proposed a structured audit methodology, there remains a significant gap in practical analysis tools for this emerging ecosystem.
-Most critically, current solutions fail to address the temporal dimension of TON security, specifically, the ability to systematically explore message ordering permutations and their impact on contract state.
-Static analysis tools such as TONScanner cannot model runtime message interleaving, while symbolic execution approaches like TSA become computationally intractable when analyzing multiple message sequences.
-In the present study, we introduce BugMagnifier, an advanced debugging instrument designed specifically for TON smart contract analysis.
-This tool complements our earlier theoretical framework by providing: (1) dynamic analysis capabilities through precise message queue manipulation, (2) enhanced vulnerability detection mechanisms for temporal vulnerabilities via message order permutations, and (3) real-time debugging features tailored to TON’s unique architecture, including full transaction phase visibility and cell-level state inspection.
-Unlike existing solutions, BugMagnifier operates at the message granularity level, enabling systematic exploration of execution paths that span multiple transactions, a capability essential for identifying the temporal vulnerabilities that dominate TON’s security landscape.
-Through this combined theoretical and practical approach, we aim to bridge the gap between Ethereum’s mature audit methodologies and the evolving needs of the TON ecosystem.
-BugMagnifier specifically addresses the critical deficiency in current tooling by providing the first practical solution for systematic race condition detection in TON smart contracts, a vulnerability class responsible for the majority of critical security incidents in the ecosystem.
+The security of smart contracts has matured into a significant research domain, though its focus remains predominantly on the Ethereum Virtual Machine (EVM) and its synchronous execution model.
+This section reviews the foundational work in smart contract vulnerability analysis, the tooling built for the EVM-centric paradigm, and the emerging research on transaction ordering in adversarial environments.
+We then contextualize existing TON-specific analysis within this broader landscape, concluding with a comparative analysis that highlights the gap filled by our work: a dynamic transaction simulator for the asynchronous TON blockchain.
+III-A General Smart Contract Vulnerabilities and Empirical Studies
+The systematic study of smart contract vulnerabilities was pioneered by researchers analyzing the EVM and Solidity.
+Early work established taxonomies of common flaws, such as reentrancy and integer overflows, which have become the primary targets for automated audit tools.
+This foundation is crucial, yet the specific vulnerability patterns are often tied to the semantics of the EVM opcodes and Solidity’s execution model.
+Recent empirical studies continue to deepen this understanding by analyzing large-scale contract deployments.
+For instance, Chen et al.
+[Chen2019b] conducted a large-scale empirical study focusing on the critical task of identifying control flow in smart contracts, a foundational step for many advanced static analyzes.
+Their work on Ethereum contracts highlights the methodologies and challenges of reasoning about contract execution paths, which differs significantly in TON’s message-passing architecture.
+Furthermore, Li et al.
+[Li2025a] provided a contemporary empirical analysis of inconsistent state update vulnerabilities.
+While their study is grounded in the EVM paradigm, the core concept–unexpected state interleaving due to concurrent operations–is directly analogous to the race conditions endemic to TON’s asynchronous model, underscoring the universal nature of state consistency challenges in decentralized systems.
+III-B Analysis Tools for the EVM and Synchronous Paradigms
+The maturity of Ethereum security research is evidenced by a robust ecosystem of analysis tools employing static, dynamic, and formal methods.
+Static analysis frameworks like Slither [Feist2019] have set a high standard; by converting Solidity source code into an intermediate representation and applying data-flow and taint analysis, they efficiently detect dozens of common vulnerability patterns without execution.
+Automated invariant generation techniques [Liu2025a] further enhance this by inferring and checking the properties that must hold across all executions, providing a higher level of assurance for synchronous contract models.
+Similarity-based analysis [Chen2025] offers another angle by comparing contract codebases to known vulnerable patterns, efficiently scaling to large ecosystems.
+Dynamic analysis and fuzzing tools, such as Echidna [Grieco2020] and xFuzz [Xue2024], take a complementary approach by executing contracts with generated or mutated inputs to uncover unexpected runtime behavior.
+Dynamic exploit generation [Wang2022c] advances this further by automatically creating concrete proofs-of-concept for detected vulnerabilities.
+On the formal verification front, tools like the Certora Prover represent the high-assurance end of the spectrum, using mathematical proofs to guarantee that a contract’s implementation satisfies its formal specification.
+Similarly, correct-by-design approaches such as VeriSolid [Nelaturu2023] enable the specification and verification of interacting contracts’ state machines, offering strong guaranties for synchronous systems like Ethereum.
+Whilst powerful, these tools are architected around the assumptions of the EVM: atomic transaction execution, a globally ordered block sequence, and synchronous state transitions.
+They reason about a single transaction’s effect on a contract’s state in isolation.
+Consequently, they lack the fundamental primitives to model TON’s core challenges: the non-deterministic interleaving of multiple independent messages targeting the same contract and the state visibility across the five distinct processing phases (storage, credit, compute, action, bounce).
+A tool like Slither cannot analyze the flow of messages between accounts, and a fuzzer like Echidna cannot systematically permute the order of a message queue to discover race conditions.
+TABLE I: Comparative Analysis of Smart Contract Security Approaches
+Tool / Approach Target Platform Core Technique Models Async. Queues? Primary Strength Limitation for TON Asynchronous Vulnerabilities
+Slither [Feist2019] EVM Static Analysis × Fast detection of code pattern vulnerabilities in Solidity. EVM-centric; cannot model TON’s message lifecycle or cell semantics.
+Echidna [Grieco2020] EVM Fuzzing × Property-based testing via execution with random inputs. Generates linear transaction sequences, not interleaved, concurrent messages.
+MEV Research [Daian2020] EVM Economic Analysis * Models profit from adversarial transaction ordering. Focuses on miner economics, not contract-state race condition detection.
+VeriSolid [Nelaturu2023] EVM Formal Verification / Model Checking × Correct-by-design construction and verification of interacting state machines.
+Designed for synchronous atomic transactions; cannot model TON’s message lifecycle or interleavings.
+Concurrency Testing [Elmas2013, Terragni2025] Generic Software Dynamic Analysis (concepts) Systematic exploration of thread/interleaving bugs.
+Framework not adapted for blockchain message-passing or TVM state.
+TONScanner [Song2025] TON Static Analysis × Detects 8+ TON-specific code defect patterns in FunC. No execution; cannot simulate message queues or detect ordering flaws.
+TSA [Espirito2025] TON Symbolic Execution × Finds runtime errors Path explosion limits exploration of multi-message permutations.
+BugMagnifier (Proposed) TON Dynamic Simulation & Permutation Testing Systematic race condition detection via controlled message orchestration. N/A (Designed to address the above limitations)
+* Analyzes ordering within a synchronous block, not an asynchronous message queue.
+III-C Adversarial Transaction Ordering and MEV Research
+Research into Maximum Extractable Value (MEV) has extensively explored the adversarial manipulation of transaction ordering within a block, a concern that shares a conceptual lineage with TON’s race conditions.
+The seminal work by Daian et al.
+[Daian2020] in “Flash Boys 2.0” documented how bots can front-run, back-run, or sandwich transactions in Ethereum’s mempool to extract profit, leading to consensus instability and economic centralization.
+This body of work formally treats blockchain state as a function of an ordered transaction list and examines the consequences of perturbing that order.
+Our work is philosophically aligned with this research in its focus on ordering-dependent state outcomes.
+However, a critical distinction exists: MEV research primarily studies the economic exploitation of ordering in a public mempool for profit within a synchronous block.
+In contrast, TON’s vulnerability is architectural: the lack of atomicity means that even two honest, correctly formatted messages can create a financially damaging race condition based solely on their arrival order, regardless of miner/validator intent.
+BugMagnifier applies the core investigative principle–systematically exploring ordering permutations–to this different, platform-inherent security problem.
+III-D Foundations in Concurrent System Verification
+The core challenge of detecting state-dependent race conditions in TON is conceptually analogous to the classic problem of verifying correctness in concurrent software, where non-deterministic thread scheduling can lead to Heisenbugs.
+Research in this domain has developed robust methodologies for systematic exploration.
+Frameworks like CONCURRIT introduced programmable thread interleaving for testing, allowing developers to specify and exhaustively explore specific scheduling scenarios of interest [Elmas2013].
+Similarly, research on differential testing of concurrent classes provides formal methods for identifying behavioral divergences between different implementations under all possible interleavings [Terragni2025].
+BugMagnifier adapts this foundational principle to the blockchain domain.
+It replaces the abstraction of thread scheduling with the concrete, controlled manipulation of TON’s message queue.
+While traditional tools reason about shared memory and locks, BugMagnifier operates on TVM state transitions and cell semantics.
+This connection positions our work not as an ad hoc TON-specific solution, but as a principled application of established concurrent system verification techniques to the novel execution model of asynchronous blockchains.
+III-E Prior Work on TON Smart Contract Security
+Security analysis for TON is a nascent field.
+Initial efforts have rightly focused on porting established techniques to the new platform.
+Song et al.’s TONScanner [Song2025] represents a critical first step, applying static analysis patterns to FunC source code to identify coding errors like unchecked bounced messages or precision loss.
+Similarly, the Ton Symbolic Analyzer (TSA) [Espirito2025] employs symbolic execution on TVM bytecode to find runtime errors.
+These tools are invaluable for detecting intra-transaction and single-message bugs, analogous to the first generation of Ethereum tools.
+For Ethereum, formal verification tools like Certora Prover represent the high-assurance end of the security spectrum, using mathematical proofs to guarantee that a contract’s implementation satisfies its formal specification [Certora2025].
+While such tools provide the strongest guaranties for synchronous blockchains, their adaptation to TON’s asynchronous model is an open challenge.
+Verifying temporal properties–such as the absence of race conditions across all possible message interleavings–requires sophisticated modeling of the message queue and TVM phases, a problem analogous to the formal verification of distributed, concurrent systems.
+BugMagnifier’s permutation testing can be viewed as a dynamic and practical stepping stone towards such formal guaranties, helping developers identify flawed contracts and gather concrete execution traces that could later inform the creation of formal models and specifications.
+Their limitation, as highlighted in our introduction, is their inability to address the inter-message and temporal vulnerabilities that define TON’s unique threat model.
+Static analysis cannot simulate the execution of a message queue, and symbolic execution becomes computationally intractable when modeling all possible permutations of even a small set of concurrent messages.
+This creates a gap where a contract can pass all existing TON linters and verifiers, yet remain highly vulnerable to race conditions.
+III-F Comparative Analysis and Identified Gap
+As illustrated in Table I, existing tools are bifurcated into two categories: 1) mature solutions for a synchronous paradigm that cannot model TON architecture, and 2) nascent TON-specific tools that cannot analyze the temporal dependencies arising from concurrency.
+Research on transaction ordering, while insightful, does not provide a practical testing framework for TON smart contracts.
+BugMagnifier is designed to bridge this gap directly.
+It takes the adversarial ordering exploration principle from MEV research and implements it as a practical, dynamic analysis tool for the TON blockchain.
+By providing granular control over the TVM’s execution phases and the message queue, it enables the systematic discovery of vulnerabilities that are invisible to all other approaches in the TON security toolkit.
+Our work thus advances the field by adapting a core security analysis methodology–permutation testing–to the unique challenges of asynchronous, sharded blockchain platforms.
 IV Proposed Solution
@@
 Our evaluation revealed critical limitations in existing tools.
-TON Contract Executor [18] abstracts away all phases except compute, bypassing essential TVM opcodes such as ACCEPT_MSG_VALUE and STORING.
+TON Contract Executor [TONContractExecutor] abstracts away all phases except compute, bypassing essential TVM opcodes such as ACCEPT_MSG_VALUE and STORING.
 This creates dangerous blind spots for vulnerabilities involving storage fees or bounce message handling.
-Similarly, TVM-linker [19] implements storage fees as fixed overhead rather than through TON’s dynamic cell depth calculations and omits bounce phase processing entirely.
+Similarly, TVM-linker [TVMLinker] implements storage fees as fixed overhead rather than through TON’s dynamic cell depth calculations and omits bounce phase processing entirely.
 Both tools fail to preserve intermediate TVM states between phase transitions, rendering them incapable of detecting state inconsistencies that occur during partial execution sequences.
-TON Sandbox [20] emerged as the only viable foundation due to its transparent implementation of TON’s transaction lifecycle.
+TON Sandbox [Sandbox] emerged as the only viable foundation due to its transparent implementation of TON’s transaction lifecycle.
 Its transaction processor module exposes the complete TVM state machine with precise gas accounting at each opcode execution, preserving full context of intermediate states between phase transitions.
@@
 The BugMagnifier interactive console is a powerful tool that provides developers with the ability to emulate TON transactions locally and identify vulnerabilities in smart contracts through advanced message management and simulation of various execution scenarios.
-The tool is publicly available on GitHub [21].
+The tool is publicly available on GitHub [BugMagnifier].
 Built as a TypeScript application leveraging Node.js runtime, BugMagnifier implements a layered architecture with distinct components for contract compilation, transaction simulation, state management, and analysis.
@@
 2[, queue <path>] [, help]
-The console provides comprehensive message queue control through commands in the table I.
-TABLE I: BugMagnifier Command Reference
+The console provides comprehensive message queue control through commands in the table II.
+TABLE II: BugMagnifier Command Reference
 Command Functionality
@@
 V Illustrative Vulnerable Smart Contract
-As part of the illustrative numerical experiment, we employ the BugMagnifier tool to demonstrate the complexities of detecting race conditions in TON smart contracts.
-To facilitate this analysis, we designed a deliberately vulnerable FunC smart contract implementing a deposit pool mechanism that exhibits critical race conditions under concurrent message processing.
-This example reveals how subtle implementation details in asynchronous blockchain environments can lead to severe financial inconsistencies, even in seemingly straightforward contracts.
-The vulnerability stems from improper state management during the ownership assignment phase, a critical weakness that becomes apparent only when analyzing message sequence permutations.
-The test contract implements a deposit pool mechanism centered around two core state variables: a 64-bit integer representing the accumulated balance (balance) and a slice storing the owner’s address (owner_address).
-Upon initial deployment, the contract establishes a zero balance and a null owner address (represented by an empty slice).
-Transaction processing logic is implemented through two primary operations: OP_ENLIST (value 1) for balance deposits and OP_CLAIM (value 2) for withdrawals.
-The critical vulnerability emerges from the contract’s ownership assignment mechanism, which triggers during the first deposit operation.
-This implementation pattern, while seemingly logical for single-user initialization, creates a fundamental race condition when multiple deposit messages arrive concurrently.
-The following listing presents the complete vulnerable contract implementation with critical sections annotated for clarity:
+We employ BugMagnifier to demonstrate the detection of race conditions inherent to TON’s asynchronous execution model.
+For this analysis, we designed a system of smart contracts implementing a deposit pool with a critical vulnerability that manifests under concurrent message processing.
+This example illustrates how the non-deterministic ordering of messages in TON can lead to severe financial inconsistencies.
+The vulnerability originates from non-atomic state management during ownership assignment, a flaw that becomes apparent only when examining permutations of messages from independently deployed user agents.
+The test system comprises a vulnerable main contract and multiple user agent contracts, modeling real-world TON applications where users interact via their own contract instances, potentially deployed across different shards.
+The core of the system is a main contract that manages a deposit pool using two primary state variables: a 64-bit integer ‘balance‘ and a slice owner_address.
+Upon deployment, the balance is zero and the owner address is empty.
+The contract processes two operations: OP_ENLIST (value 1) for deposits and OP_CLAIM (value 2) for withdrawals.
+Each user (e.g., Alice and Bob) operates through a UserContract agent, which stores the main contract’s address and forwards deposit requests.
+The vulnerability is in the ownership assignment logic within the main contract’s OP_ENLIST handler.
+Ownership is assigned to the sender of the first deposit message.
+This pattern creates a race condition when multiple user agents concurrently send their first deposit messages, as their processing order is non-deterministic in TON’s sharded architecture.
+The following listings present the vulnerable contracts.
+Main Contract (MainContract.fc):
 1#include "imports/stdlib.fc";
@@
 6(int, slice) _load_state() {
-7 slice ds = get_data()
-8 .begin_parse();
-9 int balance = slice_bits(ds) >= 64 ?
-10 ds.load_uint(64) : 0;
-11 slice owner_address = ds;
+7 slice ds = get_data().begin_parse();
+8 int balance = 0;
+9 slice owner_address = "";
+10 // Load shared state: balance and owner
+11 if (slice_bits(ds) >= 64 + 267) {
+12 balance = ds~.load_uint(64);
+13 owner_address = ds~.load_msg_addr();
+14 }
 15 return (balance, owner_address);
-13}
-14
-15(slice) get_msg_sender(cell in_msg) {
-16 slice cs = in_msg.begin_parse();
-17 int flags = cs.load_uint(4);
-18 slice sender_address =
-19 cs^.load_msg_addr();
-20 return sender_address;
-21}
-22
-23() _save_state(int balance,
-24 slice owner_address) impure {
-25 set_data(begin_cell()
-26 .store_uint(balance, 64)
-27 .store_slice(owner_address)
-28 .end_cell()
-29 );
-30}
-31
-32() recv_internal(int msg_value,
-33 cell in_msg, slice in_msg_body)
-34 impure {
-35 var (balance, owner_address) =
-36 _load_state();
-37 slice sender_address =
-38 get_msg_sender(in_msg);
-39
+16}
+17
+18() recv_internal(int msg_value, cell in_msg, slice in_msg_body) impure {
+19 var (balance, owner_address) = _load_state();
+20 slice sender_address = get_msg_sender(in_msg);
 21 if (slice_bits(in_msg_body) >= 32) {
-41 int op = in_msg_body
-42 .load_uint(32);
+22 int op = in_msg_body~.load_uint(32);
+23 // DEPOSIT OPERATION: balance increment and owner assignment
 24 if (op == OP_ENLIST) {
 25 balance += msg_value;
-45 // CRITICAL VULNERABILITY POINT
-46 // Ownership assignment occurs ONLY
-47 // during first deposit when
-48 // owner_address is empty (0 bits)
+26 // VULNERABILITY: Check-then-act on shared state ‘owner_address‘.
+27 // Two concurrent messages can both pass the empty check before either updates the state.
 28 if (slice_bits(owner_address) == 0) {
-50 _save_state(balance,
-51 sender_address);
+29 _save_state(balance, sender_address);
 30 } else {
-53 _save_state(balance,
-54 owner_address);
+31 _save_state(balance, owner_address);
 32 }
@@
 34 }
+35 // CLAIM OPERATION: withdrawal by current owner
 36 if (op == OP_CLAIM) {
-59 // Ownership verification
-60 // before withdrawal
-61 if (equal_slices(owner_address,
-62 sender_address)) {
+37 if (equal_slices(owner_address, sender_address)) {
+38 send_raw_message(begin_cell()
+39 .store_uint(0x18, 6)
+40 .store_slice(sender_address)
+41 .store_coins(balance)
+42 .store_uint(0, 1 + 4 + 4 + 64 + 32 + 1 + 1)
+43 .store_uint(OP_CLAIM, 32)
+44 .end_cell(),
+45 64
+46 );
 47 balance = 0;
-64 _save_state(balance, null());
+48 _save_state(balance, "");
 49 }
@@
 52 }
-69}
-70
-71() recv_external(slice _) {
+53}
+Listing 1: Vulnerable Main Contract with Inter-Contract Race Condition
+User Agent Contract (UserContract.fc):
+1#include "imports/stdlib.fc";
+2const OP_SEND_ENLIST = 1;
+3
+4() send_enlist(slice main_address, int amount) impure {
+5 cell msg = begin_cell()
+6 .store_uint(0, 6) // internal, no bounce
+7 .store_slice(main_address) // MainContract address
+8 .store_coins(amount) // deposit amount
+9 .store_uint(0, 1 + 4 + 4 + 64 + 32 + 1 + 1)
+10 .store_uint(OP_ENLIST, 32) // operation for MainContract
+11 .end_cell();
+12 send_raw_message(msg, 3); // mode 3: carry remaining balance
+13}
+14
+15() recv_internal(int msg_value, cell in_msg, slice in_msg_body) impure {
+16 if (slice_bits(in_msg_body) >= 32) {
+17 int op = in_msg_body~.load_uint(32);
+18 if (op == OP_SEND_ENLIST) {
+19 int amount = in_msg_body~.load_uint(128);
+20 slice main_address = _load_main_address();
+21 send_enlist(main_address, amount); // Cross-contract invocation
 22 return ();
-73}
-74
-75(int, slice) get_state() method_id {
-76 return _load_state();
-77}
-The vulnerability manifests through non-deterministic message ordering during concurrent execution, a direct consequence of TON’s asynchronous architecture.
-Consider the following scenario involving three messages:
+23 }
+24 }
+25}
+Listing 2: User Agent Contract for Inter-Shard Communication
+Consider a scenario with three contract instances:
+  • UserContract for Alice (address addr_alice)
+  • UserContract for Bob (address addr_bob)
+  • MainContract (address addr_main)
+Both Alice and Bob simultaneously initiate their first deposits via their agents, generating the following messages to the main contract:
   1. 1.
-    OP_ENLIST from Alice (sender ID 1)
+    OP_ENLIST from addr_alice
   2. 2.
-    OP_ENLIST from Bob (sender ID 2)
+    OP_ENLIST from addr_bob
   3. 3.
-    OP_CLAIM from Alice (sender ID 1).
-When these messages arrive concurrently, TON’s message queue processing creates two distinct execution paths with divergent outcomes:
-Scenario 1: Bob’s message processes first
+    OP_CLAIM from addr_alice (attempting to withdraw the pooled balance)
+Due to TON’s asynchronous architecture, the processing order of messages 1 and 2 is non-deterministic, leading to divergent financial outcomes:
+Scenario A: Message 1 (Alice) processes first.
   1. 1.
-    Bob’s OP_ENLIST executes first
+    Alice’s deposit executes. owner_address is empty, so Alice becomes the owner.
   2. 2.
-    Contract detects empty owner_address (0 bits)
+    Bob’s deposit executes. owner_address is now non-empty, so Bob is not set as owner.
   3. 3.
-    Bob’s address becomes permanent owner
+    Alice’s claim executes. Ownership verification succeeds; Alice withdraws the full balance.
   4. 4.
-    Alice’s subsequent OP_CLAIM fails ownership verification
-  5. 5.
-    Final state: balance = 20,000,000 nanotokens, owner = Bob.
-Scenario 2: Alice’s message processes first
+    Final state: Balance = 0, Owner = (empty).
+Scenario B: Message 2 (Bob) processes first.
   1. 1.
-    Alice’s OP_ENLIST executes first
+    Bob’s deposit executes. owner_address is empty, so Bob becomes the owner.
   2. 2.
-    Contract detects empty owner_address
+    Alice’s deposit executes. owner_address is now non-empty; ownership remains with Bob.
   3. 3.
-    Alice’s address becomes permanent owner
+    Alice’s claim executes. Ownership verification fails; the withdrawal is rejected.
   4. 4.
-    Alice’s OP_CLAIM succeeds (ownership verified)
-  5. 5.
-    Final state: balance = 0, owner = null.
-This race condition stems from the contract’s dependence on the slice_bits(owner_address) check during ownership assignment, a state transition that should be atomic but is instead exposed to message interleaving.
-The vulnerability is particularly insidious because:
-  • It passes standard functional tests with sequential message processing.
-  • It only manifests under specific message order permutations.
-  • The financial impact (complete loss of funds) occurs without explicit error conditions.
-  • It exploits TON’s asynchronous execution model where message processing order is non-deterministic.
-From a technical perspective, the vulnerability arises during the compute phase of transaction execution when the TVM processes the conditional ownership assignment.
-The critical instruction sequence occurs at the bytecode level as:
-1DUP // Duplicate owner_address slice
-2SLICELEN // Calculate bit length
-3PUSHINT 0 // Compare with zero
-4IFEQ // If empty, assign new owner
-This sequence creates a race window where concurrent messages can both pass the empty check before either completes state update, a classic check-then-act vulnerability amplified by TON’s message queue architecture.
-BugMagnifier enables systematic detection of this vulnerability through its message permutation capabilities.
-By generating all possible execution sequences of the three-message scenario (6 permutations), the tool identifies two distinct final states:
-  1. 1.
-    States where Alice becomes owner and withdraws funds (3 permutations)
-  2. 2.
-    States where Bob becomes owner and funds remain locked (3 permutations).
-The differential state analysis reveals critical inconsistencies in both the balance and ownership fields, with the diff command showing:
-1State difference detected:
-2- balance: 20000000 vs 0
-3- owner_address: [Bobs address] vs [empty]
-This case study exemplifies how TON’s asynchronous execution model transforms seemingly benign state transitions into critical vulnerabilities.
-The contract’s failure to implement proper synchronization mechanisms, such as nonce-based ownership claims or explicit state locking, demonstrates the heightened security requirements for asynchronous blockchain development.
-Our numerical experiments (Section VI) quantify how detection complexity varies with message ratios, revealing that such vulnerabilities become increasingly difficult to identify as message counts grow asymmetrically, a direct consequence of TON’s probabilistic message processing order.
+    Final state: Balance = (total deposits), Owner = addr_bob. Alice loses her funds.
+This is a classic check-then-act race condition.
+The check slice_bits(owner_address) == 0 and the subsequent state update are not atomic.
+Concurrent messages from different senders can interleave between these steps, causing the contract to assign ownership based on a transient, outdated state view.
+The vulnerability is subtle because:
+  • It passes linear, single-user functional testing.
+  • It manifests only under specific message orderings.
+  • It results in a complete, silent loss of funds for one party.
+  • It is a direct consequence of TON’s non-deterministic inter-shard message ordering.
+BugMagnifier systematically uncovers this flaw through permutation testing.
+For the three-message sequence, it explores all $3!=6$ possible orderings.
+The diff analysis reveals two distinct final state classes corresponding to the two scenarios above, clearly identifying the non-deterministic outcome and the critical divergence in the owner_address field.
+This case study exemplifies how asynchronous execution transforms simple state transitions into critical vulnerabilities, necessitating tools like BugMagnifier for detection.
 VI Numerical Experiments
@@
 For each configuration, For each configuration, 100 independent trials recorded the iteration count where state divergence first occurred, providing a quantitative measure of a contract’s susceptibility to race conditions.
-The experimental results are presented in Table II.
-TABLE II: Experimental Results
+The experimental results are presented in Table III.
+TABLE III: Experimental Results
 $n_{1}$ (Alice) $n_{2}$ (Bob) Total Iter. Avg. Iter. Std. Dev.
@@
 Minimum detection iterations (3.06) occur at balanced message counts ($n_{1}=n_{2}=32$), while extreme imbalances require significantly more iterations (29.32 for $n_{1}=1$, 15.46 for $n_{1}=512$).
-Higher standard deviations at imbalanced ratios indicate greater variance in the detection timing.
+Higher standard deviations at imbalanced ratios indicate a greater variance in detection timing.
 The theoretical model follows:
@@
 where $p=\frac{n_{1}}{n_{1}+n_{2}}$ and $q=\frac{n_{2}}{n_{1}+n_{2}}$ represent the relative message probabilities.
-Figure 1 plots experimental results against $\log_{2}(n_{1}/n_{2})$ with error bars representing $\pm$1 standard deviation.
+Figure 1 plots the experimental results against $\log_{2}(n_{1}/n_{2})$ with error bars representing $\pm$1 standard deviation.
 The x-axis spans from -5 (Bob dominance) to 4 (Alice dominance), with zero representing balanced messages.